OpenStudy (anonymous):

Solve 6^(2x) > 5^(x + 2).

OpenStudy (anonymous):

2xlog6>(x+2)log5 /log5 /log5 2.23x>x+2 -x -x 1.63x>2 x>1.23

OpenStudy (anonymous):

thats what i think the answer is

OpenStudy (anonymous):

the 4 answeres are x>1.83 x>1.63, x>.85, or x>.82 :(

OpenStudy (anonymous):

lol 4 answers wdh

OpenStudy (anonymous):

that doesnt make sense

OpenStudy (anonymous):

it's multiple choice :/

OpenStudy (anonymous):

so answer is x>1.63

OpenStudy (anonymous):

i see you had 1.63x, and if you divide 1.63 by 2 you get .815 which roudns up to .82

OpenStudy (anonymous):

but 1.63 and x are on the same side so nvm that

OpenStudy (anonymous):

oops yeah, this stupid windows calc sucks

OpenStudy (anonymous):

jhanny prob has right answer

OpenStudy (anonymous):

jhanny stopped typing :3

OpenStudy (anonymous):

well the answer is 1.63 i accently put that 2.23-1=1.63 but it equals 1.23

OpenStudy (anonymous):

so 2/1.23 = 1.63

OpenStudy (anonymous):

ohhhh okay

OpenStudy (anonymous):

again, stupid calc

OpenStudy (anonymous):

i shouldve used my real life one

OpenStudy (anonymous):

hehe

OpenStudy (jhannybean):

\[\large 6^{(2x)}> 5^{(x+2)}\]\[\large \ln(6)^{(2x)}> \ln5^{(x+2)}\]\[(2x)*\ln(6)> (x+2)*\ln(5)\]\[2x*\ln(6)>x*\ln(5)+2*\ln(5)\]\[2x*\ln(6)-x*\ln(5)> 2*\ln(5)\]

OpenStudy (jhannybean):

Soo...i'm needing some help.. haha.

OpenStudy (anonymous):

ewwwwwwww i need a pen for this

OpenStudy (anonymous):

Lol I think we got the right answer, jhanny :p

OpenStudy (anonymous):

wow she went nerd rage using nat log :OO, but regular log is best

OpenStudy (anonymous):

easier imo

OpenStudy (jhannybean):

Or wait... \\[x(2\ln(6)-\ln(5))>2\ln(5)\]\[\frac{ 2(\ln(6)) }{ \ln(5) }> \frac{ 2\ln(5) }{ x }\]\[\frac{ 1 }{ 2\ln(5) }*\frac{ 2\ln(6) }{ \ln(5) }> \frac{ 1 }{ x }\]

OpenStudy (jhannybean):

Yeah i'm done...... lol i don't even know what i'm doing anymore.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

yeah that is way off i think

OpenStudy (jhannybean):

what was your answer choices?

OpenStudy (anonymous):

he had i see you had 1.63x, and if you divide 1.63 by 2 you get .815 which roudns up to .82

OpenStudy (anonymous):

lol lolol

OpenStudy (anonymous):

You guys wanna give a whirl to this one?

OpenStudy (anonymous):

It's the last question in my test :3 idk how to do it

OpenStudy (anonymous):

who me?

OpenStudy (anonymous):

yessss sir. you seem to have a good head on your shoulders.

OpenStudy (anonymous):

jhanny can take this one :P ill try it tho

OpenStudy (anonymous):

not to rush, but there's 6 mins left ont he test :3

OpenStudy (jhannybean):

wait a minute,you're taking a test right now and asking us questions? Wth.

OpenStudy (anonymous):

it's online.

OpenStudy (anonymous):

unsupervised

OpenStudy (jhannybean):

That is not allowed on Openstudy. Sorry.

OpenStudy (anonymous):

wow someone is mad

OpenStudy (jhannybean):

That is called cheating, whether online OR offline.

OpenStudy (anonymous):

take it easy jhanny

OpenStudy (jhannybean):

Follow the rules of conduct for this site please.

OpenStudy (anonymous):

okok i'll just go guess. take a chill pill :3

OpenStudy (anonymous):

i think answer is 0.101

OpenStudy (anonymous):

or 0.01

OpenStudy (anonymous):

yeah i guessed that. just quick did it on my calculator.

OpenStudy (anonymous):

i mean 0.1 which is d

OpenStudy (anonymous):

final on my last exam: 76% :))

OpenStudy (jhannybean):

Alexis, can you stop giving him answers, if he's doing a test,let him finish his test, then the person can come get answers,or help.

OpenStudy (anonymous):

ok soz, btw if he cheats its his problem, i give answers to all cheaters or not

OpenStudy (anonymous):

its his choice, let him not learn or learn at will

OpenStudy (jhannybean):

But you're not going to feed him answers for a test he's currently taking, that's breaking OpenStudy rules. How can you contribute to that? I don't understand.

OpenStudy (anonymous):

well rules are rules, i did not read them and i shall not read them, but if you say so i shall stop

OpenStudy (anonymous):

well test is over and i'l never need this site again <3 it's been a good run, everybody :) can't say i'll miss doing math thanks for the help alexis!

OpenStudy (jhannybean):

Ridiculous.

OpenStudy (anonymous):

ahaha ok gl mattishot

OpenStudy (anonymous):

yeah you too :)